Exhibit 4.1 The balance sheet and income statement shown below are for Koski Inc. Note that the firm has no amortization charges, it does not lease any assets, none of its debt must be retired during the next 5 years, and the notes payable will be rolled over. Balance Sheet (Millions of $) Assets 2019 Cash and securities $4,200 Accounts receivable 17,500 Inventories 20,300 Total current assets $42,000 Net plant and equipment $28,000 Total assets $70,000 Liabilities and Equity Accounts payable $22,509 Accruals 14,391 Notes payable 6,000 Total current liabilities $42,900 Long-term bonds $11,000 Total liabilities $53,900 Common stock $3,542 Retained earnings 12,558 Total common equity $16,100 Total liabilities and equity $70,000 Income Statement (Millions of $) 2019 Net sales $105,000 Operating costs except depreciation 97,650 Depreciation 2,100 Earnings before interest and taxes (EBIT) $5,250 Less interest 1,020 Earnings before taxes (EBT) $4,230 Taxes 1,058 Net income $2,538 Other data: Shares outstanding (millions) 500.00 Common dividends (millions of $) $888.30 Int rate on notes payable & L-T bonds 6% Federal plus state income tax rate 40% Year-end stock price $60.91 ​ ​ ​ Refer to Exhibit 4.1. What is the firm's total debt to total capital ratio? Do not round your intermediate calculations. Group of answer choices 43.14% 47.76% 58.04% 53.93% 51.36%
Answer #1

TOTAL DEBT = NOTES PAYABLE + LONG TERM DEBT = 6000 + 11000 = 17000

TOTAL EQUITY = 16100

TOTAL CAPITAL = TOTAL DEBT + TOTAL EQUITY = 17000 + 16100 = 33100

DEBT TO CAPITAL RATIO = TOTAL DEBT/TOTAL CAPITAL = 17000/33100 = 0.51359 = 51.36% (rounded to 2 decimals)

ANSWER : 51.36% (Thumbs up please)